OpenStudy (anonymous):

simplify cos^2x+cos^2xtan^2x

OpenStudy (anonymous):

you can take a cos^2x out making it cos^2x(1+tan^2x) =cos^2x (sec^2x) =(cos^2x/1)(1/cos^2x) =cos^2x/cos^2x =1

OpenStudy (anonymous):

my options are 2 0 -2cot^2(theta) 2cot^2(theta)

OpenStudy (anonymous):

welp

OpenStudy (anonymous):

cos^2(x) + cos^2(x) tan^2(x) rewrite tangent^2(x) cos^2(x) + cos^2(x) [sin^2(x)/cos^2(x)] cancel the cosines cos^2(x)+[1/cos^2(x)]=cos^2(x)+sec^2(x)